### v1

# Date of publication xxxx 00, $0$, date of current version xxxx 00, $0$. n/a

Abstract
The shift from a linear to a circular economy has the potential to simultaneously reduce uncertainties of material supplies and waste generation. To date, the development of robotic and, more generally, autonomous systems have been rarely integrated into circular economy implementation strategies. In this review, we merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent physics-based research framework to lay the foundations of circular flow designs of materials, and hence, to speed-up the transition from linearity to circularity. Then, we discuss opportunities for robotics in circular economy.
Index Terms:
Automation for circularity, circular flow design, circular supply chain, thermodynamical material network.
[…]
=-21pt
I Introduction
Visualize for a moment that a built environment and human activities are contained within a plenipotentiary unit or compartment. This compartment has an input flow and an output flow of materials: the former comes from natural reserves, whereas the latter is the generation of waste and pollution. The central issue with the input flow of the compartment is that the supplies of several raw materials have a high level of risk and are also needed to manufacture high-tech and green technologies essential to a modern society [1]. Hence, these materials are labeled as “critical” by the European Union [2] and the United States [3].
The central issue with the output flow of the compartment is that environmental pollution is gradually altering the balance of ecosystems. Climate change is a manifestation of this alteration and it is mainly caused by the accumulation of greenhouse gases in the atmosphere such as carbon dioxide [4]. Other manifestations are the rise of the global temperature and the sea levels, and more frequent extreme weather events and patterns such as hurricanes or heat waves (depending on the region) [5]. In a 2024 report, the United Nations projected 2.684, 3.229, and 3.782 billion tons of waste generated in 2030, 2040, and 2050, respectively, if urgent action is not taken [6]. The Organization for Economic Co-operation and Development (OECD) reported that the share of municipal solid waste landfilled in the OECD area has decreased from an estimated 53% to 40% between 2000 and 2021, with some countries no longer using landfills (Switzerland, Germany, Denmark, Finland, Sweden, and Japan) [7].
The circular economy (CE) paradigm aims in minimizing both the input and the output flows of the plenipotentiary unit or compartment through an increase of maintenance, reuse, remanufacturing, recycling, and material efficiency [8].
Our contributions: To increase the implementation of circularity strategies at both local and global scales, in this review:
  • We merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent research framework. Our framework has a double benefit: namely, it integrates robotics into the systemic perspective needed to address circular economic challenges; and it improves the scientific foundation of the theory of CE.
  • We discuss opportunities for robotics in a circular economy to reduce the existing gap between Industry 4.0 techniques and Sustainability [9].
The rest of the paper is organized as follows. Section II covers the literature review relevant for our theoretically-coherent research framework by focusing on four different topics. Then, in Section III, we present a unification of the four key topics discussed in Section II, whereas in Section IV we collect existing models from the literature aligned with our framework. In Section V, we focus on the opportunities for robotics in a circular economy. Finally, in Section VI, we give some conclusions.
II THREE REPRESENTATIVE TOPICS FOR CIRCULAR ECONOMY
This section covers the literature regarding the four topics mentioned in the title and intersected in this review; namely, circular economy (CE), compartmental dynamical thermodynamics (CDT), robotic manipulation (RM), and deep-learning-vision (DLV).
II-A Circular Economy
Velenturf et al. [10] view circular economy as “[…] an emerging practical ideology that lacks an evidence-based theoretical framework to guide implementation […]”. Kirchherr et al. [11] identified 114 definitions of circular economy to highlight the lack of clarity and agreement about its meaning. This lack of evidence-based and clear foundations is one of the main motivations for this review. Here, we adopt the popular and intuitive definition of the Ellen MacArthur Foundation [8], which looks at closing the loops of material flows similarly to what we discussed in Section I. The large number of definitions of CE goes in parallel with the large number of indicators developed to measure circularity [12, 13, 14]. For example, De Pascale et al. [15] identified 61 indicators.
The interplay between CE and Industry 4.0 is another frequent topic in the literature. Overall, the gap between the two realms is still existing and the latter has the potential to accelerate the actualization of the former via, for example, big data, machine learning, internet-of-things, robotics, and automation [16]. In particular, Laskurain-Iturbe et al. [17] identified additive manufacturing and robotics as the most promising technologies. Kumar et al. [18] identified “high investment” and “technological immaturity” as two of the factors hindering the adoption of Industry 4.0 to enhance circularity in the food sector. Trust and common ambitions between members of CE projects were identified as very important in [19] to establish the required new collaborations between supply chain players, while the adoption of CE for small-medium enterprises was studied by Howard et al. [20].
II-B Compartmental Dynamical Thermodynamics
Thermodynamics is universal, and hence, in principle, it applies to everything in nature—from simple engineering systems to complex living organisms to our expanding universe. The laws of thermodynamics form the theoretical underpinning of virtually all disciplines in science and engineering, and hence, it is not surprising that thermodynamics can be used to model, plan, and control circular material flows [21]. The fundamental and unifying concept in the analysis of a circular economy is the concept of energy flow and material flow. The energy of a state of a dynamical system is the measure of its ability to produce changes (motion) in its own system state as well as changes in the system states of its surroundings. These changes occur as a direct consequence of the energy flow between different subsystems (or compartments) within the dynamical system.
In [22, 23, 24, 25, 26], the author merges classical thermodynamics with dynamical systems theory to develop a post-modern framework for thermodynamics predicated on open (i.e., input-state-output) interconnected dynamical systems that exchange matter and energy with the environment in accordance with the first law (conservation of energy) and the second law (non-conservation of entropy) of thermodynamics. Using a dynamical systems perspective, [26] provides a system-theoretic framework for non-equilibrium thermodynamics using a state-space formalism involving nonlinear compartmental dynamical system models characterized by energy conservation laws that are consistent with basic thermodynamic principles. This is in sharp contrast to classical thermodynamics wherein a static input-output description of the system is used. If a physical system possesses conservation properties externally, then there exists a high probability that the system also possesses these same properties internally. Specifically, if the system possesses conservation properties internally and does not violate any input-output behavior of the physical system established by experimental evidence, then the state space model is theoretically credible.
A state-space formulation is thus essential for developing a thermodynamic model with enough detail for describing the thermal behavior of heat and dynamical energy. In addition, such a model is crucial in accounting for internal system properties captured by compartmental system dynamics characterizing conservation laws, wherein subsystem energies and matter can only be transported, stored, or dissipated but not created. Furthermore, an input-state-output model is ideal in capturing system interconnections with the environment involving the exchange of matter, energy, or material, with inputs serving to capture the influence of the environment on the system, outputs serving to capture the influence of the system on the environment, and internal feedback interconnections—via specific output-to-input assignments—serving to capture interactions between subsystems. The state space model additionally enforces the fundamental property of causality, that is, nonanticipativity of the system, wherein future values of the system inputs do not influence past and present values of the system outputs. Thus, compartmental dynamical thermodynamics is ideal in providing a dynamical systems framework for modelling, planning, and controlling circular material flows in a circular economy.
II-C Robotic Manipulation for Reprocessing
Processing waste items in a circular economy perspective with automated and robotic systems presents several challenges. Kiyokawa et al. [27] identified three main elements that are needed in sorting automated systems: $i$) the end effector, which has to grasp and manipulate objects with different shapes and properties, often in a dirt and contaminated environment; $ii$) a sensing system, often based on vision, able to recognize the shape, position, category, and property of the waste item to be processed; and $iii$) a system able to plan and control stable and efficient trajectories and operation sequences.
Grasping and manipulating objects is one of the most relevant and challenging abilities required for robotic systems in several applications [28]. While traditional grasp models were based on simplified assumptions, often considering rigid objects with a limited number of contacts, approximated as points in [29], current robotic gripper developments and applications exploit more complex conditions. A rigid approach both in modeling and in the design of robotic grippers and end effectors is not suitable for getting simple, robust, effective, and reliable solutions in unstructured environments. The current trend in robotic end effector developments is towards non-rigid, or soft structures, as reviewed by Zaidi et al. [30]. Such systems are able to perform complex tasks, for example grasping and manipulating deformable objects, as the device presented by Marullo et al. [31], or multiple objects, as described by Li et al. [32]. Tactile sensing is important to detect grasp conditions and stability, especially in inherently underactuated soft grippers [33]. Soft robotic grippers can implement novel and interesting grasping strategies, for example exploiting environmental constraints, as introduced by Eppner et al. [34] and implemented by Pozzi et al. [35]. Physical constraints guiding grasping tasks can be embedded as specific features in soft robotic grippers and hands [36]. All these challenges (unstructured workspace, presence of dirt and contaminants, and deformable and fragile objects) are greatly present in systems for waste manipulation [37].
Notwithstanding the increasing need for robotic systems in waste sorting, available robotic grippers and end effectors are often not designed to cope with the large variation in material types and shapes encountered in waste material sorting, often operating with shredded material streams. As underlined by Billard and Kragic in [28], vacuum systems are the most efficient solutions when the task consists of grasping and removing an object from a set, without manipulating it; for example, in waste sorting applications. These systems are efficient and can easily adapt to different objects [38, 39, 40, 41]. In [42] an algorithm for defining the optimal suction point in plastic waste sorting is defined. Another solution able to comply with waste item variability, employing air flow, is presented by Engelen et al. [43].
When the task requires dexterity or manipulation ability, multi-fingered robotic hands, grippers, or other end effectors are needed. The development of soft robotic grippers and hands, exploiting soft materials and underactuation, has represented in the last decade an effective solution for obtaining the suitable adaptability and versatility requirements needed in unstructured environments, with simple and robust devices. This type of robotic system also represents an interesting solution in waste management systems [44]. Compliant materials and system underactuation provide adaptability to environmental uncertainties and system robustness, but lead to a new set of challenges in robot planning and control. Specific control strategies, inspired by human behavior and based on environmental constraint exploitation, have been developed in [35, 34]. Robotic grippers introduced in [45, 46] employ parts designed for smart environmental constraint exploitation [47]. One interesting, bio-inspired structure implemented in soft grippers is the Fin Ray® effect [48]. Almanzor et al. [49] developed a soft gripper based on this principle for litter sorting.
Different structures can be combined to improve gripper capabilities. For example, Sadeghian et al. [50] presented a trimodal gripper including a vacuum suction mechanism, a nano-polyurethane adhesive mechanism, and a set of claws is developed. Another interesting solution is presented in [51], where Kirigami robotic grippers are developed to manipulate different objects, including items in waste sorting systems. Bonello et al. [52] introduced an automated system for dry waste sorting and for this purpose developed a versatile gripper for waste manipulation. The design of a robotic gripper specifically developed for glass manipulation in waste sorting is presented by Kang et al. in [53].
Besides grasping and manipulating tasks, often in automated systems in the circular economy contexts involve more articulated operations. One example is represented by disassembling elements [54], where complex bimanual operations are required; for example unscrewing, removing components, cutting, etc. [55, 56, 57]. Disassembly operations are becoming particularly important in waste electronic and electrical equipment (WEEE) [58]. Autonomous systems have been proposed for disassembly tasks of specific items, e.g., electronic components [59, 60], batteries [61], cellphones [62], press fitted components [63], etc. For disassembling tasks, multifunctional grippers enabling bimanual tasks have been presented in [64, 65].
Due to their complexity, completely automated systems are still not very diffused, while robotic systems are more frequently introduced in collaboration with human operators, as reviewed by Hjorth et al. [66]. Examples of human/robot collaborative cells for disassembly tasks have been proposed in [67, 63]. Both in autonomous and collaborative disassembly tasks, proper grasp planner algorithms [68] and specific force sensing systems [69] are defined.
II-D Deep-Learning Vision for Robotics
The breakthroughs in deep learning over the last decade have significantly improved the accuracy, speed, and generalization power of vision systems. More recently, the robotics community is leveraging such advances to enhance the ability of robots to navigate and interact with environments more complex than research testbed laboratory environments. Lee et al. [70] addressed contact-rich tasks by training a deep neural network on RGB, depth, force-torque, and proprioception data. Vision-based navigation were developed in [71, 72]; the former used a deep-reinforcement-learning approach, whereas the latter combined model-based control with learning from images.
Object tracking is a useful ability for a robot that needs to interact with a dynamic environment. Object tracking systems were developed in [73, 74, 75, 76, 77, 78]. A literature review on deep-learning-based visual tracking can be found in [79]. An effective technique for visual tracking is to use object detection on each video frame, but annotating data manually is very time consuming. Hence, Horváth et al. [80] automated the annotation process. Another important task for intelligent robots is estimating the pose of target objects, especially for their manipulation. Pose estimation of objects was developed in [81, 82, 83, 84].
Visual servoing typically consists of a combination of one or more cameras, the image processing module, and the control algorithms that translate the image information and other sensor data into robot actions. Deep-learning visual servoing was proposed in [85, 86, 87]. Costante et al. [88] proposed a data-driven odometry module that provides information about the uncertainty of its estimations.
Deep-learning vision has also been proposed for autonomous marine debris detection [89, 90, 91], whereas non-marine waste was the focus of [92, 93, 94, 95]. Ku et al. [37] used a deep neural network to find an effective grasping pose for sorting construction and demolition waste. Grasping waste items is particularly challenging since their shape is difficult to predict. In contrast, manufacturing processes have less uncertainties since the geometry of the products are known. A similar consideration holds for disassembly; the conditions of an end-of-life device are difficult to predict, which makes the automation of their disassembly sequence more challenging than assembling processes in manufacturing, where the assembly sequence is known.
Jahanian et al. [96] focused on electronic waste for disassembly operations, whereas [97, 98, 99] detected the screws since their removal is the initial step of many disassembly sequences. Medical waste was considered in [100, 101]. Visual material recognition can improve both the effectiveness of robotic grasping (e.g., lower control gains with fragile items) and the automation of material sorting for a circular economy [102]. Material recognition from images was performed in [103, 104, 105].
Along with the type of material, the mass of an object is another useful property with a double application; the knowledge of the mass can make manipulation more adaptive (e.g., higher control gains with heavier items) and can also enable an autonomous quantification of material stocks to improve natural resources management in a circular economy context [106, 102] (i.e., to automate the process of material quantification in a certain area). Mass estimation from images was proposed in [107, 108, 109, 110].
III UNIFICATION OF DEEP-LEARNING VISION, COMPARTMENTAL THERMODYNAMICS, ROBOTIC MANIPULATION, AND CIRCULAR ECONOMY
In this section, we combine the four topics mentioned in the title and covered in Sections II-AII-D. The graphical intersection of the four topics is depicted in Fig. 1, where each stage of the supply and recovery chain is a thermodynamic compartment indicated by $c^{k}_{i,j}$ following the nomenclature of [21].
Figure 1: Systemic view for deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation coherently combined to design circular flows of materials. Nomenclature adopted from Zocco et al. [21].
This representation of a life-cycle is referred to as “compartmental diagram” in [21]. The origin $O$ of the frame of reference is the tail of the position vector of an infinitesimal cube of material which crosses the compartments. The compartments are connected by material flows (the arrows) and the material type is indicated by $\beta_{q}$, with $q\in[1,Q]$. The total number of compartments is $n_{c}$. The “Transportation” compartment usually does not modify the material, hence the input flow is of the same type as the output flow. The recovery chain can be automated to reduce the costs of waste sorting and re-manufacturing operations [101]. The robot can be equipped with a camera and a gripper for handling waste materials or end-of-first-use products. Then, the reprocessed materials or products can reenter the supply chain at different stages; recycled material could be used by a manufacturer to reduce its dependence on finite natural resources, while a repaired product could be transported directly to a new consumer.
Topic 1 (CE): The Rankine cycle is one of the simplest power cycles and it is usually covered in introductory courses on engineering thermodynamics (see, for example, Section 8.2 in [111]). The Rankine cycle is made of four main components (i.e., compartments); a turbine, condenser, pump, and boiler. The aim of the cycle is to generate, via a turbine, mechanical work and produce electrical energy. The cycle is closed, that is, the working fluid circulates in loop within the four compartments. If we view the working fluid as a material, then we say that the Rankine cycle is a circular life-cycle. Hence, we can think of the Rankine cycle, introduced in 1859 by the Scottish engineer William J. M. Rankine, as one of the first circular systems. The compartmental diagram in Fig. 1 for a Rankine cycle would have 8 compartments connected in a loop, 4 of which are the pipes transporting the working fluid between the 4 main components mentioned above; in addition, it would have $Q=1$ since the chemical composition of the working fluid is constant over the cycle.
Topic 2 (CDT): Along with circularity, the Rankine cycle has another important feature: its modeling begins by considering, in turn, the thermodynamics of each component (i.e., the turbine, condenser, pump, and boiler). Hence, the modeling is based on compartmental dynamical thermodynamics, which conceives, as discussed in Section II-B, multi-machine systems as a set of thermodynamic compartments connected by material flows [26]. Recall that the simplest analysis of a Rankine cycle is at steady state, i.e., when the rate of change of energy $\dot{E}=0$. Under the simplifying, but practically useful, assumption of steady state, the term “dynamical” in “compartmental dynamical thermodynamics” is removed since, at equilibrium, the dynamics of the system is not considered.
Topic 3 (RM): As demonstrated in [101], the mechanics governing the dynamics of a robot can be derived from an energy balance. Hence, industrial manipulators can be considered as thermodynamic compartments. Recalling the analogy with the Rankine cycle, an industrial manipulator performing waste sorting is a thermodynamic compartment analogous to the turbine or the pump in the Rankine cycle. Now, given that a robotic manipulator is a thermodynamic compartment, we can assert that robotic manipulation significantly affects the performance of the robot-compartment, and hence, the performance of the network of compartments as a whole. Therefore, an effective manipulation is essential for an effective recovery chain as depicted in Fig. 1.
Topic 4 (DLV): Consider again a manipulator as a thermodynamic compartment. If it is equipped with one or more cameras as in Fig. 1, then these deep-learning-vision algorithms can process the images in real-time to improve the ability of the robot-compartment to handle waste materials or end-of-first-use products.
IV A BIBLIOTHECA OF COMPARTMENTS
Since mass and energy balances are general natural principles, they have been used to develop many dynamical models to depict the dynamics of mechanical, chemical, and biological systems. In the context of this paper, these existing models are a library of thermodynamic compartments that can be added and connected as appropriate to design networks of thermodynamic compartments as the one depicted in Fig. 1. They can also work as a starting point for developing custom compartments for a particular material flow design. Existing models of thermodynamic compartments are discussed below.
In [101] (specifically, in Proposition 1) the standard form of robot dynamics is derived from an energy balance, and thus, a manipulator can be considered a thermodynamic compartment. The robot manipulator can be used for waste sorting or product disassembly in reprocessing stages as shown in Fig. 1. Similarly, a wheeled vehicle can be modeled as a thermodynamic compartment since its dynamics can be derived from the first law of thermodynamics (see [112] and Proposition 1 in [101]). In Fig. 1, for the case where materials are moved on a truck (the case of a large vehicle) or if a robot is used to collect litter (the case of a small vehicle) [49], a wheeled vehicle can be thought as a transportation compartment.
Whenever the waste is located in areas with difficult access, an aerial drone may be a preferred mode for a transportation compartment [113]. The Lagrangian mechanics for an aerial drone is given in [114]. The large amount of marine debris is becoming of increasing concern since microplastics can be ingested by fishes and, eventually, by humans [115]. To remove the debris materials and re-circulate them into production chains, underwater vehicles can be used [90]. Hence, in this case, one of the transportation compartments in Fig. 1 would be an underwater vehicle, which recovers the waste materials from the ocean (the environment is indicated by $c^{n_{c}}_{n_{c},n_{c}}$). The Lagrangian dynamics for an underwater vehicle is given in [116].
Thus far in this section, we covered four types of robot compartments, namely, manipulators, wheeled vehicles, aerial drones, and underwater vehicles. Their dynamics can be derived from energy or power balance equations (i.e., from the dynamical form of the first law of thermodynamics). There are of course other types of thermodynamic compartments whose models may include one or more mass balances.
Compared to fossil fuels, bio-fuels are more carbon neutral because they are sourced from plantations, are more renewable, and can be produced locally to make a community self-sustained. Hence, bio-fuels are more aligned with circular economy principles than their fossil counterpart [117]. A network of thermodynamic compartments for the biomass storage and production of bio-methane was designed in [21], where the transportation-compartment is a vehicle whose dynamics was covered above, while the model of the anaerobic digester follows from mass balances [118, 119].
The reuse of products or materials requires a preliminary washing step if they are infected. This is the case for devices used in healthcare such as surgical scissors, syringes, and inhalers. Hence, a circular flow of infected materials could involve a washing compartment along with transportation and robotic waste sorting. The thermodynamic model of a dishwasher heater was developed in [120].
The study of material flows often involves the comparison between circular and linear systems, which are both shown in Fig. 1; the former considers one or more material flow loops, whereas the latter sends material straight to incineration or the environment (indicated by $c^{n_{c}}_{n_{c},n_{c}}$). For these comparative studies, the dynamics of an incinerator compartment is useful and it is provided in [121] as a result of both mass and energy balances.
V OPPORTUNITIES FOR ROBOTICS IN CIRCULAR ECONOMY
Opportunities in application of robot technologies in a circular economy are wide-ranging satisfy two key aspects connected with waste management: $i$) substitution of human activities in the case of elevated health risk due to the nature of the process or to the waste material properties; and $ii$) performance increment in waste sorting outputs or waste collection quality, avoiding high costs due to destination of improper material to the wrong waste treatment process in the new prospective of a circular economy. Waste separated collection [122] and waste sorting were widely investigated with automation technologies focusing on the high level of waste flows to be managed and optimizing the high margins with respect to domestic waste management. The promising prospects of reducing management costs and increasing the effectiveness and efficiency of selection processes has stimulated the in-depth study of innovative robotization technologies [123], in the face of a notable quantity of industrial automation solutions on a market scale with sustainable solutions applied to municipal solid waste [124, 125].
The transition from linear waste management to a circular economy requires a notable leap in quality in the performance of all phases of the process, in order to avoid effects of economic inefficiency due to the multiplication of transport and management costs due to flows affected by unsuitable materials with respect to the quality conditions required by recycling processes. From this perspective, robotization technologies can play a decisive role in completion of automatic management of supply chains, in the final classification of flows destined for recycling in the face of reduced levels of unwanted or contaminated materials, replacing processes that are still largely based on manual selection [126].
Further promising prospects for the application of robot technologies are: $i$) the disassembly processes of machinery and electronic equipment in general; and $ii$) the management of processes with high health risk. The first category of activities, which includes a notable number of research and development initiatives using robot technologies, is undoubtedly of considerable interest for the prospect of extending the concepts of circular economy to high economical value waste flows reducing the amount of precious materials today sent to disposal. The second sector of activity appears less investigated, but it is clear that interesting prospects (even if for waste flows of limited scope) can be applied to processes for the classification or identification of highly dangerous or radioactive materials erroneously present in transport processes or storage. In this case, robotization of on-site processes could reduce intervention times and the use of highly specialized personnel with notable benefits in all areas of logistics, which today are seriously damaged by the possible occurrence of incorrect deliveries in light of the controls now present in entrance to treatment plants. Similar application of on-site robotic processes are experienced and developed for the management of demolition and construction waste flows [94, 127].
VI CONCLUSION
In this paper, we coherently combined the literature related to four key topics: deep-learning vision, compartmental dynamical thermodynamics, robotic manipulation, and circular economy. Specifically, the latter is the area of application wherein the other three topics are presented.
Deep-learning vision is advancing rapidly due to performance improvements of neural networks, but its use for the automation of recovery tasks such as waste sorting and disassembly remains minimal. The adaptability and scalability of deep-learning vision can enable highly-flexible waste sorting facilities to process very different types of materials and products. This is currently not feasible with traditional automated methods, which can process only particular classes of items. In addition, deep-learning vision can be used for wide-area autonomous materials monitoring to improve natural resources management (as detailed in [106]).
Another key ingredient for highly-flexible waste sorting systems are advanced robotic manipulators capable of handling items with shapes difficult to predict such as those of end-of-life products. However, advanced manipulation such as the one based on soft components poses new challenges for planning and control of robot motions. To align the grippers with circular economy principles, the use of bio- or bio-degradable materials in their designs is an option worth exploring. The benefits of an effective autonomous waste sorting system include a reduction in contamination risks by human operators, a reduction in operating costs, and an increase of the rate of material separation for reuse. An inefficient separation causes a higher fraction of material being sent for incineration or, illegally or not, to the environment, thus becoming litter and marine debris.
The integration of deep-learning vision and robotic manipulation can also be used in disassembly operations to repair faulty items. Given that, to date, the technological level in manufacturing is significantly more advanced than in re-manufacturing, there are significant opportunities in transferring the knowledge from the former to the latter. This will certainly be a challenging task since the geometries and conditions of end-of-life products are much less predictable than the ones of new items. Disassembly for the recovery of precious metals and critical raw materials is particularly promising due to the economic value of these natural resources, which justifies the costs of robotic systems.
The term “circular” itself implies a systemic perspective that considers the material life-cycle from extraction to its reuse and disposal. Hence, this review proposes compartmental dynamical thermodynamics to establish rigorous physics-based theoretical foundations of circular flow designs. Systems such as robots, transportation systems, and incinerators, obey the laws of thermodynamics, and hence, their interactions form a network of thermodynamic compartments. The thermodynamic compartments can be added, modified, or removed at the design stage of the compartmental network in order to achieve a circular material flow. As we showed in this review, a theoretically-coherent physics-based framework underpinning the design of circular flow systems is missing in the literature and it is needed to speed-up the implementation of circularity at the local and global scales.
ACKNOWLEDGMENT
F. Zocco sincerely thanks the Department of Information Engineering and Mathematics of University of Siena for hosting him for this research.